OpenStudy (anonymous):

Im too lazy to do this. What is the simplest form for: -265/1000

OpenStudy (anonymous):

\[-\frac{ 53 }{ 200 }\]

jimthompson5910 (jim_thompson5910):

Step 1) Identify the GCF of 265 and 1000. The GCF is 5 Step 2) Divide each part by the GCF. 265/5 = 53 1000/5 = 200 So that is why -265/1000 = -53/200

OpenStudy (anonymous):

- Thanks , another question.. This is my homework and Im too lazy to do dis. Whats: -1 3/10

OpenStudy (anonymous):

Really hope people are too lazy to answer

OpenStudy (anonymous):

You mean this? \[-(1 + \frac{ 3 }{ 10 })\] It's simply.. \[-\frac{ 13 }{ 10 }\]

OpenStudy (anonymous):

-No. Like Negative One and 3/10.. @Light&Happiness